OpenStudy (anonymous):

if f(x)=3x+9, g(x)=5*2 find (f+g)(x)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (debbieg):

(f+g)(x) simply means to add together the expressions for f(x) and g(x) (and simplify, is possible). E.g. (f+g)(x)=f(x)+g(x)

OpenStudy (debbieg):

I'll do an example similar to yours: if f(x)=4x-7, g(x)=6*8 find (f+g)(x) ANSWER: (f+g)(x)=f(x)+g(x)=4x-7+6*8 =4x-7+48 =4x+41